# ROADWork: A Dataset and Benchmark for Learning to Recognize, Observe, Analyze and Drive Through Work Zones

### v2

@@ C.1.1 Other Interesting Recognition Scenarios
Figure 15: Adapting to New Geographies. (a) Starting from Pittsburgh, we progressively add data from new cities to train a detector, leading to significant accuracy gains (+7.4 $AP$ with all cities). (b) Geographic adaptation remains challenging. Training on source data only serves as our baseline, while training on both source and target data represents the upper bound. Adaptation methods such as 2PCNet [43] provide limited improvement over the baseline. For example, the upper-bound gap of adaptation method for “barricade” (-13.8 $AP_{50}$) and for “vertical panel” (-20.7 $AP_{50}$) is very large.
Adapting to New Geographies. While we discovered work zones in new geographies (Figure 5 of the main manuscript), does our recognition model maintain the same performance? Domain adaptation methods have explored geographic adaptation, but mainly across countries [101, 94, 42, 114, 35] and mostly for common objects like cars [101]. Obtaining supervised data for new geographies , such as new cities in our case, is expensive to scale. We make two observations: (a) A geographic domain gap exists in our data, and (b) state-of-the-art adaptation methods do not address this gap.
@@
Following state-of-the-art adaptation methods [43, 114], we use a Faster R-CNN Resnet50 backbone for training, assuming different levels of available target domain data. Training on source data only is our baseline (red in Figure 15 (b)), whereas upper bound is trained on both labeled source and labeled target data (depicted in green). State-of-the-art adaptation methods [43, 114] (depicted in blue) do not significantly improve adaptation over baseline. Compared to the upper bound in Figure 15 (b), performance gaps remain for heavily represented objects like cones (-5.4 $AP_{50}$) and drums (-12.6 $AP_{50}$), and also for rare objects like barricades (-13.8 $AP_{50}$) and vertical panels (-20.7 $AP_{50}$). Our ROADWork dataset highlights the geographic domain gap problem, underscoring the need for new algorithms to bridge this gap.
Method $AP$ $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
@@
Table 16: Label Unification. We train a bounding box detector on unified Cityscapes [17] and ROADWork label space by (a) pseudo-labeling our dataset using a pretrained Cityscapes [17] (CS) model (b) via UniDetector [102]. Testing on Cityscapes [17] and compared to a pretrained model solely trained on Cityscapes [17] labels, we observe significant degradation when trained on unified label space while UniDetector [102] improves the performance over naive pseudo-labeling.
Label Unification. Suppose we aim to train a unified detector that detects both common objects like cars from a common driving dataset and rare objects from the ROADWork dataset simultaneously. This requires label unification. While practical, unifying the label space is a challenging task [113, 102] – training a model on a unified set of categories reduces performance compared to specialized models individually trained on each dataset. One reason is due to the presence of unlabeled instances of a particular category in the unified dataset, another could be due concept overlaps between two labels in the unified dataset. To assess this observation, we consider the Cityscapes [17] bounding box dataset in addition to our ROADWork dataset, using UniDetector [102] with a Faster R-CNN model. We train our model on the unified label space of Cityscapes (common objects) and ROADWork dataset (long-tailed objects) – (a) by employing a pre-trained detector (see Section 3) to pseudo-label the ROADWork dataset. (b) employing the method proposed by UniDetector [102]. When testing our unified models on the Cityscapes validation set, we observe a considerable drop in performance (-5.0 AP) when naive pseudo-labeling is used. However, UniDetector [102] closes the gap with the Cityscapes pre-trained model by improving the performance by +2.0 $AP$ over naive pseudo-labeling.
@@ C.2 Analyzing Work Zones
LLaVA-1.5 [60] 7B 0.4 11.0 9.4 0
LLaVA-1.5 [60] 13B 0.3 9.8 8.0 0
LLaVA-NEXT [60, 50] 13B 0.2 9.4 6.9 0
LLaVA-NEXT [60, 50] 34B 0.3 9.3 6.9 0
@@
LLaVA-1.5 [60] 7B 27.0 24.7 48.0 112.1
LLaVA-1.5 [60] 13B 27.7 25.1 48.6 113.1
LLaVA-NEXT [50] 13B 28.2 25.3 48.3 116.4
LLaVA-NEXT [50] 34B 28.4 26.1 47.2 113.2
Table 17: Newer and Larger Vision-Language Models (VLMs). Consistent with the poor performance trends in Section 5 of the main manuscript, larger pretrained VLMs (13B–34B parameters) also fail to describe work zones. Switching to a newer generation of VLMs [50] does not improve performance, reinforcing the under-representation of work zones in existing large-scale training datasets. Fine-tuning helps, but even a 34B model provides only a marginal improvement (+1.4 METEOR).
Newer and Larger Vision-Language Models (VLMs). We performed our experiments in Section 5 of the main manuscript using LLaVA-1.5-7B. However, two key questions arise: (a) Do larger VLMs also struggle with work zones descriptions? (b) Are newer VLMs such as LLaVA-NEXT [50] better at describing work zones than older models?
As shown in Table 17, pre-trained VLMs of all sizes perform poorly on work zones unless they are fine-tuned on ROADWork dataset. Unfortunately, even the larger LLaVA-NEXT-34B model provides only a marginally performance gain (e.g. +1.4 METEOR) over LLaVA-1.5-7B [60]. Moreover, newer VLMs like LLaVA-NEXT [50] do not significantly outperform the previous generation of VLMs, likely because they still lack exposure to work zone images. Our ROADWork dataset fills that gap, advancing high-level scene understanding in work zones.
@@
Table 18: Does fine-tuning a vision-language foundation model on ROADWork cause overfitting? Large-scale vision-language models trained on millions of images risk overfit and catastrophic forgetting of prior learned distributions when trained on small target datasets. We test this by evaluating LLaVA-7B [60] models on COCO-Captions [10]. We evaluate using the pretrained model and model fine-tuned on ROADWork data. Surprisingly, the fine-tuned model does not degrade in performance and instead shows significant improvements (+18.6 ROUGE).
Does fine-tuning a vision-language foundation model on ROADWork cause overfitting? In Section C.1 we asked if fine-tuned open vocabulary models forget previously learned distributions when trained on our ROADWork dataset. A similar question applies to vision-language foundation models, which we investigate here. We evaluate two models on COCO-Captions [10], a dataset that provides captions for many real-world images. To test our hypothesis, we evaluate two models (a) a pretrained LLaVA-1.5-7B [60] model (b) a LLaVA-1.5-7B additionally fine-tuned on the ROADWork dataset. The input prompt to both the model is “Describe the given image in detail.” Our findings are reported in Table 18. Surprisingly, contrary to our expectations, the fine-tuned model performs better on almost all metrics. Further analysis suggests that the captioning style of COCO-Captions [10] validation set is more similar to the terse and direct scene descriptions of the ROADWork dataset, whereas the original LLaVA-1.5-7B [60] training data is more detailed and flowery. We hypothesize that fine-tuning on our ROADWork dataset improved model alignment for captioning tasks. We leave further investigation to future work.
@@
YNet [63] w/ Pretrained Segm. [17] 31.28 102.7 28.28 95.92 29.84 102.39 40.76 113.38
YNet [63] w/ ROADWork Segm. 22.68 80.78 22.41 75.33 21.82 83.28 30.21 84.58
Table 19: Pathway Prediction in Images. We employ YNet [63] with a segmentation model trained on Cityscapes [17] as our baseline, and train a segmentation model with ROADWork dataset, and we observe that work zone object segmentations improve pathway and goal predictions. Displacement Error (ADE) and Final Displacement Error (FDE) captures the error of predicted pathway and goal from the ground truth pathway and goal respectively. We also report results for different thresholds of average curvatures, hypothesizing that it is more difficult to navigate workzones where pathways are more irregular. We do observe that displacement errors of both predicted pathway and goal is higher at the higher curvature threshold.
@@ C.3 Driving through Work Zones
Metric $AE\%<\theta$ vs Pixel level metrics [63].  [63] presents results on pixel level metrics like Average Displacement Error (ADE) and Final Displacement Error (FDE), we also report those metrics. However, we believe $AE\%<\theta$ measures model performance more fairly in autonomous driving situations. This is because pixel level metrics like average displacement error [63] but do not account for camera’s field of view in the ego-car’s viewpoint. We have access to the camera instrinsics $K$ and the angular error is computed by finding the angle between ground truth point $p$ and predicted point $\hat{p}$ in pixel coordinates,
$AE(p,\hat{p})=\cos^{-1}\left(\frac{(K^{-1}p)\cdot(K^{-1}\hat{p})}{\left\lVert K^{-1}p\right\rVert\left\lVert K^{-1}\hat{p}\right\rVert}\right)$
Now, we define $AE\%<\theta$ as the percentage of predictions whose angular error is within a threshold $\theta$. Do note horizontal field of view of our images are around 50°50°50  °.
Pathway Prediction Results with Pixel Level metrics. Nevertheless, like [63], we also report pixel level displacement metrics. Table 19 shows goal and pathway results, Final Displacement Error (FDE) is the pixel error between the predicted goal and the actual goal while Average Displacement Error (ADE) is the error between predicted pathway and actual pathway. We observe that ROADWork improves both FDE (-21.3%) and ADE (-27.5%). We also bin the pathways in terms of curvature, and observe that paths with higher curvature are difficult to predict, however, model trained on ROADWork dataset improves FDE (-25.4%) and ADE (-25.9%) in those cases.
@@ Appendix D Implementation Details
Detecting Work Zone Objects. We employ the pre-trained open vocabulary models [105, 115, 61] as is and follow their custom vocabulary protocol. For training Mask R-CNN [33], we use the mmdetection [8] library initialized with COCO [57] weights. We use the default model zoo parameters with the 1x schedule. For DINO [110], Mask DINO [51] and DiffusionDet [9], we use their official codebases and weights. We employ the simple copy-paste implementation from mmdetection [8] and use the default parameters.
@@
Adapting to New Geographies We follow the same pre-train and adaptation protocol described in 2PCNet [43].
@@
Generating Work Zone Descriptions. To circumvent memory constraints, we train models via low rank adaptation (LORA) [36]. We also hypothesized utilzing object predictions as context would improve description quality. We compose the coarse vocabulary work zone object detector (from Section 3 in the main paper) to align our descriptions. We employ rank $R=128$ and alpha $\alpha=256$ while performing LORA fine-tuning on LLaVA-7B [59] for 4 epochs. We keep the rest of the parameters as is, following LLaVA’s training schedule. The model prompt to generate descriptions is “You are the planner of an autonomous vehicle, ONLY describe the workzone in the scene identifying and describing the spatial relationship of relevant objects to plan and navigate a route”. While training with additional object context, we use ground truth to append a programmatic prompt for each object – “(object_category: confidence) at [(x1, y1), (x2, y2)]. While testing with additional object context, we use detector predictions.
@@
Figure 16: Pathway Prediction for Work Zone Image Sequences. We show examples of trajectory heatmaps predicted by YNet [63] for video sequences. Input to the model is the image and observed pathway, also shown is the future pathway (computed from actual driving). Frames are outlined indicating plausible pathway heatmap and colliding pathway heatmap. (Sequence 1) Following vehicles is a learned cue. (Sequence 2) Exploiting available free space is also learned. (Sequence 3) Even if the initial goal is plausible, model predicts an unsafe trajectory that would collide with work zone objects. Later, model course-corrects the trajectory when closer to work zone objects.
@@
Figure 17: Pathway Prediction in Work Zone Images. We show examples of goal and trajectory heatmaps predicted by YNet [63]. Input to the model is the image and observed pathway, also shown is the future pathway (both computed from actual driving videos). Top row shows the predicted goal heatmaps while the bottom row shows the predicted pathway heatmaps, conditioned on a sampled goal. We observe that the predicted goal heatmap (marked with an orange box for clarity) is close to the ground truth goal, and the predicted pathway is plausible.
@@
Figure 18: Pathway Prediction in Work Zone Images: Failure cases. We show examples of goal and trajectory heatmaps predicted by YNet [63] where the model fails. Input to the model is the image and observed pathway, also shown is the future pathway (both computed from actual driving). Top row shows the predicted goal heatmaps (marked with an orange box for clarity) while the bottom row shows the predicted pathway heatmaps, conditioned on a sampled goal. We observe, (a-b) the model selects vehicles in front as goal without considering global semantics. (c) Modelling multimodality of goals is a challenge, model is unable to predict all goals at an intersection. (d) Even if the goal is valid, the pathway prediction fails for heavily blocked work zones.

### v3

@@ C.1.1 Other Interesting Recognition Scenarios
Figure 15: Adapting to New Geographies. (a) Starting from Pittsburgh, we progressively add data from new cities to train a detector, leading to significant accuracy gains (+7.4 APAP with all cities). (b) Geographic adaptation remains challenging. Training on source data only serves as our baseline, while training on both source and target data represents the upper bound. Adaptation methods such as 2PCNet [kennerley20232pcnet] provide limited improvement over the baseline. For example, the upper-bound gap of adaptation method for “barricade” (-13.8 $AP_{50}$) and for “vertical panel” (-20.7 $AP_{50}$) is very large.
Adapting to New Geographies. While we discovered work zones in new geographies (Figure 5 of the main manuscript), does our recognition model maintain the same performance? Domain adaptation methods have explored geographic adaptation, but mainly across countries [wang2020train, varma2019idd, kalluri2023geonet, zheng2024instance, hoyer2022daformer] and mostly for common objects like cars [wang2020train]. Obtaining supervised data for new geographies , such as new cities in our case, is expensive to scale. We make two observations: (a) A geographic domain gap exists in our data, and (b) state-of-the-art adaptation methods do not address this gap.
@@
Following state-of-the-art adaptation methods [kennerley20232pcnet, zheng2024instance], we use a Faster R-CNN Resnet50 backbone for training, assuming different levels of available target domain data. Training on source data only is our baseline (red in Figure 15 (b)), whereas upper bound is trained on both labeled source and labeled target data (depicted in green). State-of-the-art adaptation methods [kennerley20232pcnet, zheng2024instance] (depicted in blue) do not significantly improve adaptation over baseline. Compared to the upper bound in Figure 15 (b), performance gaps remain for heavily represented objects like cones (-5.4 $AP_{50}$) and drums (-12.6 $AP_{50}$), and also for rare objects like barricades (-13.8 $AP_{50}$) and vertical panels (-20.7 $AP_{50}$). Our ROADWork dataset highlights the geographic domain gap problem, underscoring the need for new algorithms to bridge this gap.
Method APAP $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$
@@
Table 16: Label Unification. We train a bounding box detector on unified Cityscapes [cordts2016cityscapes] and ROADWork label space by (a) pseudo-labeling our dataset using a pretrained Cityscapes [cordts2016cityscapes] (CS) model (b) via UniDetector [wang2023detecting]. Testing on Cityscapes [cordts2016cityscapes] and compared to a pretrained model solely trained on Cityscapes [cordts2016cityscapes] labels, we observe significant degradation when trained on unified label space while UniDetector [wang2023detecting] improves the performance over naive pseudo-labeling.
Label Unification. Suppose we aim to train a unified detector that detects both common objects like cars from a common driving dataset and rare objects from the ROADWork dataset simultaneously. This requires label unification. While practical, unifying the label space is a challenging task [zhao2020object, wang2023detecting] – training a model on a unified set of categories reduces performance compared to specialized models individually trained on each dataset. One reason is due to the presence of unlabeled instances of a particular category in the unified dataset, another could be due concept overlaps between two labels in the unified dataset. To assess this observation, we consider the Cityscapes [cordts2016cityscapes] bounding box dataset in addition to our ROADWork dataset, using UniDetector [wang2023detecting] with a Faster R-CNN model. We train our model on the unified label space of Cityscapes (common objects) and ROADWork dataset (long-tailed objects) – (a) by employing a pre-trained detector (see Section 3) to pseudo-label the ROADWork dataset. (b) employing the method proposed by UniDetector [wang2023detecting]. When testing our unified models on the Cityscapes validation set, we observe a considerable drop in performance (-5.0 AP) when naive pseudo-labeling is used. However, UniDetector [wang2023detecting] closes the gap with the Cityscapes pre-trained model by improving the performance by +2.0 $AP$ over naive pseudo-labeling.
@@ C.2 Analyzing Work Zones
LLaVA-1.5 [liu2024improved] 7B 0.4 11.0 9.4 0
LLaVA-1.5 [liu2024improved] 13B 0.3 9.8 8.0 0
LLaVA-NEXT [liu2024improved, li2024llavanext-strong] 13B 0.2 9.4 6.9 0
LLaVA-NEXT [liu2024improved, li2024llavanext-strong] 34B 0.3 9.3 6.9 0
@@
LLaVA-1.5 [liu2024improved] 7B 27.0 24.7 48.0 112.1
LLaVA-1.5 [liu2024improved] 13B 27.7 25.1 48.6 113.1
LLaVA-NEXT [li2024llavanext-strong] 13B 28.2 25.3 48.3 116.4
LLaVA-NEXT [li2024llavanext-strong] 34B 28.4 26.1 47.2 113.2
Table 17: Newer and Larger Vision-Language Models (VLMs). Consistent with the poor performance trends in Section 5 of the main manuscript, larger pretrained VLMs (13B–34B parameters) also fail to describe work zones. Switching to a newer generation of VLMs [li2024llavanext-strong] does not improve performance, reinforcing the under-representation of work zones in existing large-scale training datasets. Fine-tuning helps, but even a 34B model provides only a marginal improvement (+1.4 METEOR).
Newer and Larger Vision-Language Models (VLMs). We performed our experiments in Section 5 of the main manuscript using LLaVA-1.5-7B. However, two key questions arise: (a) Do larger VLMs also struggle with work zones descriptions? (b) Are newer VLMs such as LLaVA-NEXT [li2024llavanext-strong] better at describing work zones than older models?
As shown in Table 17, pre-trained VLMs of all sizes perform poorly on work zones unless they are fine-tuned on ROADWork dataset. Unfortunately, even the larger LLaVA-NEXT-34B model provides only a marginally performance gain (e.g. +1.4 METEOR) over LLaVA-1.5-7B [liu2024improved]. Moreover, newer VLMs like LLaVA-NEXT [li2024llavanext-strong] do not significantly outperform the previous generation of VLMs, likely because they still lack exposure to work zone images. Our ROADWork dataset fills that gap, advancing high-level scene understanding in work zones.
@@
Table 18: Does fine-tuning a vision-language foundation model on ROADWork cause overfitting? Large-scale vision-language models trained on millions of images risk overfit and catastrophic forgetting of prior learned distributions when trained on small target datasets. We test this by evaluating LLaVA-7B [liu2024improved] models on COCO-Captions [chen2015microsoft]. We evaluate using the pretrained model and model fine-tuned on ROADWork data. Surprisingly, the fine-tuned model does not degrade in performance and instead shows significant improvements (+18.6 ROUGE).
Does fine-tuning a vision-language foundation model on ROADWork cause overfitting? In Section C.1 we asked if fine-tuned open vocabulary models forget previously learned distributions when trained on our ROADWork dataset. A similar question applies to vision-language foundation models, which we investigate here. We evaluate two models on COCO-Captions [chen2015microsoft], a dataset that provides captions for many real-world images. To test our hypothesis, we evaluate two models (a) a pretrained LLaVA-1.5-7B [liu2024improved] model (b) a LLaVA-1.5-7B additionally fine-tuned on the ROADWork dataset. The input prompt to both the model is “Describe the given image in detail.” Our findings are reported in Table 18. Surprisingly, contrary to our expectations, the fine-tuned model performs better on almost all metrics. Further analysis suggests that the captioning style of COCO-Captions [chen2015microsoft] validation set is more similar to the terse and direct scene descriptions of the ROADWork dataset, whereas the original LLaVA-1.5-7B [liu2024improved] training data is more detailed and flowery. We hypothesize that fine-tuning on our ROADWork dataset improved model alignment for captioning tasks. We leave further investigation to future work.
@@
YNet [mangalam2021goals] w/ Pretrained Segm. [cordts2016cityscapes] 31.28 102.7 28.28 95.92 29.84 102.39 40.76 113.38
YNet [mangalam2021goals] w/ ROADWork Segm. 22.68 80.78 22.41 75.33 21.82 83.28 30.21 84.58
Table 19: Pathway Prediction in Images. We employ YNet [mangalam2021goals] with a segmentation model trained on Cityscapes [cordts2016cityscapes] as our baseline, and train a segmentation model with ROADWork dataset, and we observe that work zone object segmentations improve pathway and goal predictions. Displacement Error (ADE) and Final Displacement Error (FDE) captures the error of predicted pathway and goal from the ground truth pathway and goal respectively. We also report results for different thresholds of average curvatures, hypothesizing that it is more difficult to navigate workzones where pathways are more irregular. We do observe that displacement errors of both predicted pathway and goal is higher at the higher curvature threshold.
@@ C.3 Driving through Work Zones
Metric $AE\%<\theta$ vs Pixel level metrics [mangalam2021goals].  [mangalam2021goals] presents results on pixel level metrics like Average Displacement Error (ADE) and Final Displacement Error (FDE), we also report those metrics. However, we believe $AE\%<\theta$ measures model performance more fairly in autonomous driving situations. This is because pixel level metrics like average displacement error [mangalam2021goals] but do not account for camera’s field of view in the ego-car’s viewpoint. We have access to the camera instrinsics $K$ and the angular error is computed by finding the angle between ground truth point $p$ and predicted point $\hat{p}$ in pixel coordinates,
[…]
Now, we define $AE\%<\theta$ as the percentage of predictions whose angular error is within a threshold $\theta$. Do note horizontal field of view of our images are around 50°.
Pathway Prediction Results with Pixel Level metrics. Nevertheless, like [mangalam2021goals], we also report pixel level displacement metrics. Table 19 shows goal and pathway results, Final Displacement Error (FDE) is the pixel error between the predicted goal and the actual goal while Average Displacement Error (ADE) is the error between predicted pathway and actual pathway. We observe that ROADWork improves both FDE (-21.3%) and ADE (-27.5%). We also bin the pathways in terms of curvature, and observe that paths with higher curvature are difficult to predict, however, model trained on ROADWork dataset improves FDE (-25.4%) and ADE (-25.9%) in those cases.
@@ Appendix D Implementation Details
Detecting Work Zone Objects. We employ the pre-trained open vocabulary models [xu2023open, zhou2022detecting, liu2024grounding] as is and follow their custom vocabulary protocol. For training Mask R-CNN [he2017mask], we use the mmdetection [mmdetection] library initialized with COCO [lin2014microsoft] weights. We use the default model zoo parameters with the 1x schedule. For DINO [zhangdino], Mask DINO [li2023maskdino] and DiffusionDet [chen2023diffusiondet], we use their official codebases and weights. We employ the simple copy-paste implementation from mmdetection [mmdetection] and use the default parameters.
@@
Adapting to New Geographies We follow the same pre-train and adaptation protocol described in 2PCNet [kennerley20232pcnet].
@@
Generating Work Zone Descriptions. To circumvent memory constraints, we train models via low rank adaptation (LORA) [hu2022lora]. We also hypothesized utilzing object predictions as context would improve description quality. We compose the coarse vocabulary work zone object detector (from Section 3 in the main paper) to align our descriptions. We employ rank $R=128$ and alpha $\alpha=256$ while performing LORA fine-tuning on LLaVA-7B [liu2023llava] for 4 epochs. We keep the rest of the parameters as is, following LLaVA’s training schedule. The model prompt to generate descriptions is “You are the planner of an autonomous vehicle, ONLY describe the workzone in the scene identifying and describing the spatial relationship of relevant objects to plan and navigate a route”. While training with additional object context, we use ground truth to append a programmatic prompt for each object – “(object_category: confidence) at [(x1, y1), (x2, y2)]. While testing with additional object context, we use detector predictions.
@@
Figure 16: Pathway Prediction for Work Zone Image Sequences. We show examples of trajectory heatmaps predicted by YNet [mangalam2021goals] for video sequences. Input to the model is the image and observed pathway, also shown is the future pathway (computed from actual driving). Frames are outlined indicating plausible pathway heatmap and colliding pathway heatmap. (Sequence 1) Following vehicles is a learned cue. (Sequence 2) Exploiting available free space is also learned. (Sequence 3) Even if the initial goal is plausible, model predicts an unsafe trajectory that would collide with work zone objects. Later, model course-corrects the trajectory when closer to work zone objects.
@@
Figure 17: Pathway Prediction in Work Zone Images. We show examples of goal and trajectory heatmaps predicted by YNet [mangalam2021goals]. Input to the model is the image and observed pathway, also shown is the future pathway (both computed from actual driving videos). Top row shows the predicted goal heatmaps while the bottom row shows the predicted pathway heatmaps, conditioned on a sampled goal. We observe that the predicted goal heatmap (marked with an orange box for clarity) is close to the ground truth goal, and the predicted pathway is plausible.
@@
Figure 18: Pathway Prediction in Work Zone Images: Failure cases. We show examples of goal and trajectory heatmaps predicted by YNet [mangalam2021goals] where the model fails. Input to the model is the image and observed pathway, also shown is the future pathway (both computed from actual driving). Top row shows the predicted goal heatmaps (marked with an orange box for clarity) while the bottom row shows the predicted pathway heatmaps, conditioned on a sampled goal. We observe, (a-b) the model selects vehicles in front as goal without considering global semantics. (c) Modelling multimodality of goals is a challenge, model is unable to predict all goals at an intersection. (d) Even if the goal is valid, the pathway prediction fails for heavily blocked work zones.